OpenStudy (anonymous):

Evaluate the expression for a = –1 and b = 5. (ab)^2

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (-1 * 5)^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow -5^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow 25 }\)

OpenStudy (anmolsingh):

(ab)^2=(-1*5)^2=(-5)^2=25